# MS-0001-1922.65

### v1

@@ 5.1 Policy Impact: Costs
Orthopedic (OR) 15.3 349,325.3 \$ 4,723,951
Cardiovascular (CV) 29.3 17,865.0 \$ 1,810,240
Gastroenterology Urology (GU) 39.5 1,432,739.3 837,178
Hematology (HE) 50.0 147,978.9 \$ 462,295
General Hospital (HO) 42.4 141,564.6 \$ 222,750
Ophthalmic (OP) 15.8 39,984.7 46,843.2
@@ 5.2 Post-Hoc Analysis
In this section, we delve deeper into assessing the performance of our proposed policy via post-hoc analyses. We start by examining the characteristics of the deferred medical devices based on some of the top predictors of recall risk. The results are summarized in Table 8. In this table, we also compare devices that are deferred by our proposed policy for more in-depth evaluation to those that have been correctly accepted or rejected. It is evident that the risk estimates are higher for unrecalled devices deferred by our policy compared to those that have been correctly accepted. Similarly, the risk estimates are lower for recalled devices deferred by the policy compared to those that have been correctly rejected. Additionally, significant differences can be observed in variables such as Predicate Average Age, Predicate Newest Age, Number of Recalls, Weighted Number of Recalls, and Variance of Recalls between the deferred devices and the correctly diagnosed ones. These results confirm our hypothesis that while there are some easy cases suitable for direct algorithmic decision-making, some other cases are difficult and require more in-depth evaluation and human expert judgment.
Our post-hoc analysis also provide various other practice-relevant insights. For example, they highlight that while the predicted risk is highly effective in distinguishing between easy and hard cases, the risk estimate alone may not be sufficient for diagnosing the safety of deferred devices. Specifically, our results indicate that deferred devices with higher values for Num. of Recalls and Weighted Num. of Recalls are more likely to be recalled. This aligns with the recent protocols suggested by the FDA (FDA 2023b) in selecting predicates that continue to perform safely, and highlights that not only the number of recalls matters for selected predicates but also the timing of their recalls.
@@
cell12 = c=3, cell16 = c=3, hline1,4,12 = -, hline2 = 2-8, hline3 = 2,4,6,8, & Unrecalled Devices Recalled Devices
@@
Risk Estimate 0.09 (0.04) 0.29 (0.07) 0.67 (0.13) 0.30 (0.07)
Num. of Unmatched Specialties 0.18 (0.44) 0.14 (0.39) 0.11 (0.37) 0.16 (0.39)
Predicate Average Age 5.93 (5.28) 5.36 (4.84) 4.95 (4.07) 5.06 (4.56)
Predicate Newest Age 4.34 (5.21) 3.72 (4.55) 3.05 (3.31) 3.24 (4.01)
Predicate Oldest Age 7.74 (6.62) 7.21 (6.48) 7.29 (6.27) 7.32 (6.80)
Num. of Recalls 0.15 (1.29) 0.18 (0.68) 2.30 (3.10) 0.48 (3.33)
Weighted Num. of Recalls 0.08 (0.79) 1.30 (3.18) 1.66 (2.23) 0.28 (1.83)
Variance of Recalls 0.01 (0.15) 0.01 (0.04) 0.18 (0.58) 0.02 (0.19)
Table 9 presents a comparison of devices accepted by our proposed policy and those accepted based on the FDA’s current practice in the testing set. As our policy does not directly diagnose the deferred devices, this analysis is conducted under the assumption that the deferred devices are evaluated following the FDA’s current practice. We observe from our results that the risk estimate for devices accepted by our proposed policy is lower compared to devices accepted by the FDA’s current practice. The lower risk estimate indicates that our policy is cautious about accepting devices with a potentially higher risk. In addition, the age of the latest-approved predicate for devices accepted under our policy is slightly higher compared to devices accepted under the FDA’s current practice, while the age of the earliest-approved predicate is slightly lower. This indicates a careful balance between leveraging proven safety records and embracing new technology. Furthermore, the number of recalls, weighted number of recalls, and variance of predicates’ recalls for devices accepted by our policy is significantly lower than those accepted by FDA’s current practice. This is a substantial difference and indicates that our policy has the tendency to accept devices with predicates with fewer historical recalls, which aligns with the existing literature suggesting the correlation between the chance of recall and the number of recalls for predicates. Finally, we observe that the number of unmatched specialties and the oldest age of predicates do not exhibit a significant difference between devices accepted by our policy and the current practice.
@@
hline1,3,11 = -, hline2 = 2,4, & Our Policy Current Practice
@@
Risk Estimate 0.23 (0.11) 0.29 (0.19)
Num. of Unmatched Specialties 0.15 (0.40) 0.15 (0.40)
Predicate Average Age 5.53 (4.98) 5.52 (4.91)
Predicate Newest Age 3.89 (4.76) 3.79 (4.62)
Predicate Oldest Age 7.40 (6.57) 7.50 (6.62)
Num. of Recalls 0.19 (1.29) 0.51 (2.40)
Weighted Num. of Recalls 0.10 (0.72) 0.32 (1.45)
Variance of Recalls 0.01 (0.14) 0.05 (1.11)
6 Policy and Managerial Implications
@@
Our work investigates the degree to which a data-driven policy can assist the FDA in improving its 510(k) medical device clearance process. We hypothesized that our combined human-algorithm approach to evaluating medical devices will result in a reduction of recall rates and the workload burden imposed on the FDA. Conducting an in-depth evaluation of the performance of our policy as well as the FDA’s current practice, we found that while the predicted risk is highly effective in distinguishing between easy and hard cases, utilizing it alone may not be sufficient. This confirms our hypothesis that there is a need for a combined human-algorithm approach, where devices with a mid-range predicted risk of recall (non-easy cases) are deferred to human experts for further evaluation. That is, integrating the FDA’s expertise with quantitative evidence is required to improve the 510(k) medical device clearance process. A conservative evaluation of our proposed policy based on our data showed a 38.9% improvement in the recall rate and a 43.0% reduction in the FDA’s workload. Our cost analyses projected that implementing our policy could result in significant annual cost-savings ranging between \$2.4 billion and \$2.7 billion. Overall, these findings imply that our proposed policy is effective in significantly reducing recall rates, the FDA’s workload, and costs incurred due to recalls.
@@
We believe that our work addresses some of the main concerns about the current 510(k) process. Most recently, the FDA has responded to these concerns with modernization efforts “to improve the safety of medical devices while continuing to create more efficient pathways to bring critical devices to patients” (FDA 2018b). In 2018, the FDA proposed “promoting innovation and improving safety by driving innovators toward reliance on more modern predicate devices.” That is, the newer devices should be compared to the benefits and risks of more modern technology as the as older predicates might not reflect the advanced technology embedded in new devices or align with the FDA’s current understanding of device benefits and risks (FDA 2018c). However, the FDA ultimately acknowledged that its initial proposal “may not optimally promote safer and more effective devices” in all instances since devices that use modern rapidly-evolving technology could benefit from comparison to more recent predicates whereas older devices may establish a history of safety and effective use in other cases. We found that a higher value of the newest predicate age is linked to a reduced recall risk. Specifically, devices with a predicate age of less than five years are associated with a significantly higher risk of recall, possibly due to their limited real-world usage that may reveal potential issues over time. Regarding the oldest predicate age, our results indicated that both high and low values can predict a high recall risk. This aligns with discussions in the literature regarding the benefits and risks of using older predicates. Older predicates may signify safety and be a standard for patient care, but they might not reflect the latest technology, leading to compatibility issues (FDA 2018b, FDA 2018c). We also found that lower values of average and median predicate age generally correlate with a lower recall risk.
@@
Our research also addresses another of the FDA’s recent inquiries on the use of AI/ML. In 2023, the FDA published a discussion paper seeking feedback on the opportunities and challenges of using of AI/ML in the development of drugs and medical devices intended to be used with drugs (FDA 2023f). Our research addresses the crucial considerations relevant to model development, performance, monitoring, and validation within the context of utilizing AI/ML to enhance the evaluation of medical devices.
@@
One of the specific benefits of our proposed policy is that it has a transparent structure (two phases based on clear rules), which allows for continued transparency of the 510(k) review process. Second, it benefits from explainability, which is an important element in allowing stakeholders, such as regulators and health providers, to sanity check models beyond mere performance (Amann et al. 2020, Zech et al. 2018, Saghafian and Hopp 2019). Third, in developing and evaluating our proposed policy, we paid specific attention to selecting appropriate metrics, such as false positive and false native rates and the FDA’s workload. We highlight that relying solely on one of these metrics may not provide a complete picture of the reality. For example, false positives and false negatives can have significantly different consequences. False positives may lead to unnecessary rejection of safe devices, while false negatives could result in missed diagnoses of unsafe devices.
@@
Finally, our proposed policy follows findings from the recent literature that suggest combining human intuition and judgment with the power of AI/ML algorithms by creating human-algorithm “centaurs” can go a long way (Orfanoudaki et al. 2022, Saghafian 2023). Our proposed policy creates a combined human-algorithm approach by deferring some decisions to human exerts and others to a well-trained algorithm. Our results show that this approach can be very effective, mainly because it makes use of both the expertise of FDA committees when needed and the power of an algorithm that can accurately predict future recall risks for a variety (but not all) of devices.
@@
It is worth discussing possible concerns regarding the practicality of implementing our proposed policy, including (i) FDA’s ability to leverage our data-driven clearance policy, and (ii) reaction of applicants to the 510(k) pathway. Regarding point (i), we note that the FDA’s recent draft guidance on best practices for selecting a predicate device indicates that our policy would align with the FDA’s avowed commitment “to improve the predictability, consistency and transparency” of the 510(k) pathway while not proposing changes to applicable statutory and regulatory standards, such as how the FDA evaluates substantial equivalence, or the applicable requirements, including the requirement for valid scientific evidence” (FDA 2018b). As discussed earlier, three of the FDA’s proposed best practices are an attempt to improve the safety of medical devices by mitigating the use of predicates with safety issues. Our data-driven policy uses established AI/ML methods to provide additional scientific evidence, namely an estimated recall risk of a 510(k) applicant device based, in part, on safety issues present in predicate devices. Moreover, our clearance policy is not intended to be a substitute for the FDA’s expertise in assessing applicant devices. It is a tool that the FDA can leverage, as it sees fit, to increase device safety while reducing workload. The FDA, among other approaches, can choose to use our tool as a guide for identifying devices that warrant higher priority concerns rather than outright rejection.
Regarding point (ii), we note that similar to the FDA, applicants will have the capability to replicate our model. This proactive response can, in turn, decrease the workload on the FDA during the review process and reduce the risk of recalls after a device is cleared. If, on the other hand, applicants begin to only mask factors that increase the risk of rejection (e.g., changing the country of manufacture with no other substantive change), our ML predictor will adapt accordingly as the masking becomes more prominent in the training data over time. Furthermore, as discussed earlier, our clearance policy is neither intended to be a substitute for the FDA’s expertise in assessing applicant devices, nor is it intended to be a blackbox. We expect that the FDA will continue its commitment to transparency when communicating the reasons for a rejection, including the prediction and variable importance in our model. Moreover, FDA rejections of 510(k) applicants are specific to the application under review. Applicants are further able to seek clearance by submitting a new application that addresses the FDA’s concerns.
7 Conclusion
This research aims to enhance safety and expedite clearance procedures within the 510(k) pathway. Our primary focus is on introducing a data-driven clearance policy intended to assist the FDA in refining the 510(k) medical device clearance process. Our methodology and approach are informed by comprehensive discussions with our collaborator, who possesses substantial experience in various FDA regulatory consulting projects. Our modeling framework enables the FDA to integrate its expertise with quantitative evidence. While it does not prescribe a specific course of action for devices that warrant further evaluation, it allows for the incorporation of the FDA experts’ judgment when necessary. Focusing the FDA’s expertise on devices requiring the most attention can significantly enhance the evaluation process, improving patient safety and reducing unnecessary workload for the FDA.
Our results suggest that that our methodology can lead to significant potential improvements in the recall rates and the FDA’s workload. Despite this, our projections indicate a relatively low percentage of acceptance of unsafe devices. Our analysis also suggests that using our policy can lead to a substantial reduction in costs, alongside a decrease in adverse event outcomes and an enhancement in patient safety.
@@ 7 Conclusion
We believe future studies are required to further investigate the impacts of implementing our proposed policy. For example, as discussed earlier, we only provide conservative estimates on the cost-savings due to our proposed policy. Specifically, our assessment of the replacement costs for recalled devices likely underestimates the true cost of a recall by excluding the potential cost of adverse events. An expansion on our current work can assess the amount of injury and death that could have been averted under our proposed policy. Finally, we believe a future study could run an experiment, where 510(k) applicant devices are randomly assigned to either our proposed policy or the current FDA procedure. Data collected on such a randomized experiment can further inform the FDA about the advantages and disadvantages of our policy, and shed light on needed modifications prior to full implementation.

### v2

@@ 5.1 Policy Impact: Costs
Orthopedic (OR) 17.9 179,229.3 \$ 2,276,838
General Hospital (HO) 32.3 855,095.4 \$ 1,076,394
Anesthesiology (AN) 51.0 603,998.7 \$ 990,606
Cardiovascular (CV) 29.1 8,824.2 \$ 959,569
5.2 A Comparison of Human-Algorithm and ML-Only Policies
In this section, we compare the performance of our policy, which is a combined human-algorithm approach, with the performance of an ML-only policy that is solely based on non-human decisions.
The ML-only policy is a one-threshold policy obtained by solving the following optimization problem:
$\min_{t\geq 0}$ $\;\;\lambda\;\mathbb{P}(\delta^{+}\leq t)+(1-\lambda)\;\mathbb{P}(\delta^{-}\geq t),$
where $\lambda\in(0,1)$ can be interpreted as the same $\lambda$ used in our optimization problem (1)–(5). The resulting threshold $t$ is indeed the same as choosing a threshold that strikes a trade-off between sensitivity and $1-\text{specificity}$ in the ROC curve, which is a common practice.
Unlike the ML-only policy, which does not result in deferring any applicant devices, our policy results in deferring some proportion of applicant devices to the FDA committee for more in-depth evaluation in a setting with additional a priori information about the risk of the applicant device and lower workload. Consequently, the performance of our policy partly depends on the performance of the FDA committee to evaluate the deferred devices. In our setting, a conservative comparison of the ML-only policy and our policy offers no additional insight. In particular, for any risk thresholds $(\ell,h)$ selected by our policy, there exists a corresponding threshold for the ML-only policy by setting $t=h$. With this construction, both policies yield the same acceptance rate for unsafe devices and the same rejection rate for safe devices. Devices with estimated risk at or above $h$ are rejected under both policies, as the ML-only threshold is the same as the rejection boundary of our policy. Devices with risk below $\ell$ are accepted by both approaches since the single threshold $h$ is equal to or greater than $\ell$. For devices with risk between $\ell$ and $h$, the ML-only policy accepts the device, whereas our policy defers the decision to the FDA. However, because our dataset contains only devices that were cleared (accepted) by the FDA, all deferred devices will be accepted as well. Accordingly, we consider the added benefits of our policy in the comparison, including the provision of additional information for deferred devices and the reduction of the FDA’s workload. For example, they help determine the review depth and the level of scrutiny required for each applicant device. Riskier devices may undergo more rigorous evaluations, involving higher scrutiny and a more thorough analysis of safety and efficacy. Additionally, reducing the FDA’s workload allows the committees to focus on complex applicant devices that require more in-depth evaluation.
To compare the performance of our policy with the ML-only policy, we simulate the performance of the FDA to evaluate the deferred devices in the new setting. We assume that the additional information for deferred devices does not hurt the FDA committees’ performance in evaluating a deferred device. This assumption implies that all the safe devices deferred will be accepted by the FDA’s committees. Furthermore, we assume that the probability that the FDA’s committees will fail to detect an unsafe deferred device is as follows:
$\mathcal{L}(f(X),k)=2\left(1-\frac{1}{1+\exp\left(-k(f(X)-\hat{\ell})\right)}\right),$
where $f(X)$ is the estimated risk, $\hat{\ell}$ is the low threshold obtained by our policy, and parameter $k\geq 0$ is a scalar where higher values of $k$ correspond to improved performance of the FDA’s committees in detecting unsafe devices. When $k=0$, this probability is equal to $1$ by our design for any unsafe device that has been deferred. This implies that $k=0$ is the baseline that matches the FDA’s current practice in evaluating deferred devices without supplementary information.
Figure 11 illustrates the predicted risk and the probability of failing to reject an unsafe device for 100 random samples, where samples are ordered based on their predicted risk. As can be seen, the probability of failing to reject an unsafe device decreases as the predicted risk of the deferred device increases. When $k=0$, the FDA’s committees will fail to reject an unsafe device with a probability of 1. As $k$ increases, this probability decreases. This observation aligns with our intuition that unsafe devices appearing less risky are more challenging for the FDA’s committees to detect.
Figure 11: Predicted risk and probability of failing to reject unsafe deferred devices
We compare the performance of our policy with that of an ML-only policy in terms of (i) rejection of safe devices, (ii) acceptance of unsafe devices, and (iii) the percentage improvement in recall rate relative to the FDA’s current practice. We consider three values of the parameter $k\in\{L,M,H\}$ and evaluate various combinations of modeling parameters $\lambda\in(0,1)$, $\xi^{ru}\in[0.1,0.7]$ , and $\xi^{as}\in[0.1,0.7]$. The parameter $\rho$ is fixed at $0.6$, consistent with the representative policy discussed in Section 5. Figure 14 shows the Pareto frontiers for policies that reject fewer than 15% of safe devices to reflect more feasible cases in terms of implementation. As shown in Figure 14, our policy consistently outperforms the ML-only policy across all levels of $k$. The performance gap is largest when the overall rejection rate of safe devices is relatively low. We observe that the performance of our policy improves with higher levels of $k$. In other words, when the FDA committee becomes more effective at rejecting deferred unsafe devices, our policy’s partial deferral strategy becomes increasingly beneficial. In practical terms, the ML-only policy is most effective when its risk estimations are highly accurate, where the ML-only policy can make accurate decisions for hard cases. However, in many real-world scenarios, including ours, risk estimation models perform reasonably well but lack uniform accuracy across all cases. In such settings, selectively deferring difficult cases to human experts can significantly enhance decision-making and lead to a better-performing policy overall.
Figure 12: Accept unsafe vs reject safe
Figure 13: Recall rate pct. improvement vs reject safe
Figure 14: Comparison of the performance of our policy vs the ML-only policy in the testing set
@@ 5.3 Post-Hoc Analysis
In this section, we delve deeper into assessing the performance of our proposed policy via post-hoc analyses. We start by examining the characteristics of the deferred medical devices based on some of the top predictors of recall risk. The results are summarized in Table 6. In this table, we also compare devices that are deferred by our proposed policy for more in-depth evaluation to those that have been correctly accepted or rejected. It is evident that the risk estimates are higher for unrecalled devices deferred by our policy compared to those that have been correctly accepted. Similarly, the risk estimates are lower for recalled devices deferred by the policy compared to those that have been correctly rejected. For unrecalled devices, deferred cases have modestly elevated recall histories and slightly newer predicates compared to accepted ones. For recalled devices, deferred cases exhibit lower recall histories and slightly older predicates than those that were confidently rejected. These patterns suggest that deferred devices fall into a borderline zone where risk signals and related features raise some concern, but not strongly enough to support a confident accept or reject decision. This supports the need for additional human review and the design of our policy to deffer hard cases for further evaluation.
@@
cell12 = c=3, cell16 = c=3, hline1,4,11 = -, hline2 = 2-8, hline3 = 2,4,6,8, & Unrecalled Devices Recalled Devices
@@
Risk Estimate 0.04 (0.01) 0.10 (0.03) 0.32 (0.14) 0.10 (0.03)
Predicate Median Age 4.74 (4.20) 3.81 (3.33) 3.49 (2.84) 3.86 (3.44)
Predicate Newest Age 4.57 (5.26) 3.63 (4.50) 2.84 (3.64) 3.46 (4.31)
Predicate Oldest Age 7.86 (6.70) 7.08 (6.32) 7.23 (6.55) 7.41 (6.86)
Num. of Class 2 Recalls 0.07 (0.56) 0.26 (1.00) 2.84 (5.04) 0.44 (2.99)
Weighted Num. of Recalls 0.02 (0.10) 0.09 (0.23) 0.57 (0.38) 0.11 (0.26)
Variance of Recalls 0.00 (0.05) 0.01 (0.10) 0.39 (2.87) 0.02 (0.22)
Table 7 presents a comparison of devices accepted by our proposed policy and those accepted based on the FDA’s current practice in the testing set. As our policy does not directly diagnose the deferred devices, this analysis is conducted under the assumption that the deferred devices are evaluated following the FDA’s current practice. We observe that the risk estimate for devices accepted by our proposed policy is relatively similar to that of the FDA’s accepted devices (0.08 vs. 0.10). This indicates that both approaches accept devices with comparably low estimated risk; however, the predicate characteristics differs. The age of the latest-approved predicate for devices accepted under our policy is slightly higher compared to devices accepted under the FDA’s current practice, while the age of the earliest-approved predicate is relatively the same. This shows a balance between leveraging proven safety records and embracing new technology. Furthermore, the number of class 2 recalls, weighted number of recalls, and variance of predicates’ recalls for devices accepted by our policy is significantly lower than those accepted by FDA’s current practice. This is a substantial difference and indicates that our policy has the tendency to accept devices with predicates with fewer historical recalls, which aligns with the existing literature suggesting the correlation between the chance of recall and the number of recalls for predicates.
@@
hline1,3,10 = -, hline2 = 2,4, & Our Policy Current Practice
@@
Risk Estimate 0.08 (0.04) 0.10 (0.09)
Predicate Median Age 4.11 (3.66) 4.04 (3.58)
Predicate Newest Age 3.92 (4.76) 3.78 (4.65)
Predicate Oldest Age 7.35 (6.48) 7.33 (6.48)
Num. of Class 2 Recalls 0.21 (1.16) 0.49 (2.57)
Weighted Num. of Recalls 0.07 (0.20) 0.12 (0.28)
Variance of Recalls 0.01 (0.10) 0.06 (1.42)
6 Managerial Insights and Conclusion
@@
Our work investigates the degree to which a data-driven policy can assist the FDA in improving its 510(k) medical device clearance process. We hypothesized that our combined human-algorithm approach to evaluating medical devices will result in a reduction of recall rates and the workload burden imposed on the FDA. We conducted an in-depth evaluation of the performance of our policy as well as the FDA’s current practice. The results confirm our hypothesis that there is a need for a combined human-algorithm approach, where devices with a mid-range predicted risk of recall (non-easy cases) are deferred to human experts for further evaluation. That is, integrating the FDA’s expertise with quantitative evidence is required to improve the 510(k) medical device clearance process. A conservative evaluation of our proposed policy based on our data showed a 32.9% improvement in the recall rate and a 40.5% reduction in the FDA’s workload. Our cost analyses projected that implementing our policy could result in significant annual cost savings of \$1.7 billion. Overall, these findings imply that our proposed policy is effective in significantly reducing recall rates, the FDA’s workload, and costs incurred due to recalls.
@@
We believe that our work addresses some of the main concerns about the current 510(k) process. Most recently, the FDA has responded to these concerns with modernization efforts “to improve the safety of medical devices while continuing to create more efficient pathways to bring critical devices to patients” (FDA 2018c). In 2018, the FDA proposed “promoting innovation and improving safety by driving innovators toward reliance on more modern predicate devices.” That is, the newer devices should be compared to the benefits and risks of more modern technology as the as older predicates might not reflect the advanced technology embedded in new devices or align with the FDA’s current understanding of device benefits and risks (FDA 2018d). However, the FDA ultimately acknowledged that its initial proposal “may not optimally promote safer and more effective devices” in all instances since devices that use modern rapidly-evolving technology could benefit from comparison to more recent predicates whereas older devices may establish a history of safety and effective use in other cases. We found that a higher value for the newest predicate age is linked to a reduced recall risk. This is possibly due to the limited real-world usage of newer predicates, which may not have been sufficient to reveal potential issues. Regarding the oldest predicate age, our results indicated that higher values can lead to a lower recall risk. This aligns with with the idea that older predicates can indicate safety, as they have a history of safe and effective use and may be the gold standard for patient care. We also found that higher values of the median predicate age correlate with a lower recall risk. Altogether, these findings highlight a learning curve effect. That is, if a predicate device has been on the market only briefly, latent problems may remain undiscovered, leading to a higher likelihood of future recalls.
@@
One of the specific benefits of our proposed policy is that it has a transparent structure (two phases based on clear rules), which allows for continued transparency of the 510(k) review process. Second, it benefits from explainability, which is an important element in allowing stakeholders, such as regulators and health providers, to sanity check models beyond mere performance (Amann et al. 2020, Zech et al. 2018, Saghafian and Hopp 2019). Third, in developing and evaluating our proposed policy, we paid specific attention to selecting appropriate metrics, such as false positive and false negative rates and the FDA’s workload. We highlight that relying solely on one of these metrics may not provide a complete picture of the reality. For example, false positives and false negatives can have significantly different consequences. False positives may lead to unnecessary rejection of safe devices, while false negatives could result in missed diagnoses of unsafe devices.
@@
Finally, our proposed policy follows findings from the recent literature that suggest combining human intuition and judgment with the power of AI/ML algorithms by creating human-algorithm “centaurs” can go a long way (Orfanoudaki et al. 2022, Saghafian 2023). Our proposed policy creates a combined human-algorithm approach by deferring some decisions to human experts and others to a well-trained algorithm. Our results show that this approach can be effective in our setting, mainly because it makes use of both the expertise of FDA committees when needed and the power of an algorithm that can accurately predict future recall risks for a variety (but not all) of devices. To keep the human-algorithm partnership improving, one useful approach is to establish an oversight committee, separate from the review committees, that meets periodically to audit recent approvals, rejections, and deferrals, hear stakeholder feedback, and examine extreme cases flagged by our policy. The committee can then adjust the decision thresholds as needed which allows the system to improve continuously over time.
@@
It is worth discussing possible concerns regarding the practicality of implementing our proposed policy, including (i) FDA’s ability to leverage our data-driven clearance policy, and (ii) transparency vs gaming. Regarding point (i), we note that the FDA’s recent draft guidance on best practices for selecting a predicate device indicates that our policy would align with the FDA’s avowed commitment “to improve the predictability, consistency and transparency” of the 510(k) pathway while not proposing changes to applicable statutory and regulatory standards, such as how the FDA evaluates substantial equivalence, or the applicable requirements, including the requirement for valid scientific evidence” (FDA 2018c). As discussed earlier, three of the FDA’s proposed best practices are an attempt to improve the safety of medical devices by mitigating the use of predicates with safety issues. Our data-driven policy uses established AI/ML methods to provide additional scientific evidence, namely an estimated recall risk of a 510(k) applicant device based, in part, on safety issues present in predicate devices. Moreover, our clearance policy is not intended to be a substitute for the FDA’s expertise in assessing applicant devices. It is a tool that the FDA can leverage, as it sees fit, to increase device safety while reducing workload. The FDA, among other approaches, can choose to use our tool as a guide for identifying devices that warrant higher priority concerns rather than outright rejection.
Regarding point (ii), we note that some level of transparency is desirable as it steers manufacturers toward inherently safer designs and decreases the review process. However, complete disclosure of every model feature or weight may invite gaming (Wang et al. 2023). For example, applicants might begin to mask factors that increase the risk of rejection. A practical compromise is to publish the core risk factors while keeping the finer details of the prediction model and the risk thresholds of the policy proprietary to the FDA. The policy itself will be adapted dynamically, which allows detecting and counteracting some masking tactics. It is worth noting that our estimates of reductions in recall events and FDA’s workload rest on the empirical distribution of 510(k) submissions observed in our dataset. If our proposed policy alters manufacturers’ strategic behavior, for example by deterring some higher-risk devices or attracting additional low-risk devices, the post-implementation submission mix could differ from our estimation sample. This could introduce endogeneity and bias the projected gains. Periodic recalibration of the policy may help the algorithm adapt to emerging patterns and mitigate drift arising from manufacturers’ adaptive behavior, but it may not fully eliminate such effects.
Beyond FDA, our human-algorithm framework can assist other domains that sift large applicant pools under tight resource constraints. For example, quantitative screens are already used in university admissions (grade-point average, test scores, course rigor) before human committees review borderline files. A data-driven first pass like ours could formalize that practice by recommending acceptance or rejection while flagging edge cases for further deliberation.
@@
We believe future studies are needed to further investigate the impacts of implementing our proposed policy, as well as potential changes to the FDA’s existing procedures. For example, the predicate selection process that is currently determined by the applicant could be replaced or supplemented with an algorithmically identified set based on the similarity of the applicant device and the existing 510(k) devices. Finally, we believe a future study could run an experiment, where 510(k) applicant devices are randomly assigned to either our proposed policy or the current FDA procedure. Data collected on such a randomized experiment can further inform the FDA about the performance of our policy, and shed light on needed modifications prior to full implementation.